Adam is 5 years younger than Eve. In 1 year, Eve will be three times as old as Adam was 4 years ago. Find their ages now.
Answer by drk(1908)

Adam is 5 years younger than Eve
translates to
(i) A = E - 5
In 1 year, Eve will be three times as old as Adam was 4 years ago
translates to
(ii) E +1 = 3(A-4)
by substitution, we get
(iii) E + 1 = 3(E-5-4)
E + 1 = 3E - 27
2E = 28
E = 14
A = 9